# MicroAdam: Accurate Adaptive Optimization with Low Space Overhead and Provable Convergence

### v1

@@ D.3 Analysis Under PL Condition
$\mathbb{E}[f(\theta_{T+1})]-f^{*}\leq\bigl{(}1-\frac{\eta\mu}{C_{0}}\bigr{)}^{T}(f(\theta_{1})-f^{*})+\eta\left(\frac{LC_{0}\sigma^{2}+LC_{0}(C_{1}+C_{2}^{2})G^{2}}{\mu\epsilon}+\frac{(1+C_{1})G^{2}d+C_{1}G^{2}}{\sqrt{\epsilon}}\right)$
$\hfill+\,\eta^{2}\left(\frac{3L^{2}C_{0}C^{2}_{1}G^{2}}{2\mu\epsilon^{3/2}}+\frac{(1+2C_{1})C_{1}LG^{2}d}{\epsilon}+\frac{LC_{1}^{2}G^{2}}{2\epsilon}\right).$
@@
D.4 Non-convex Analysis with Weight Decay
@@
8:     $\theta_{t+1}=(1-\eta_{t}\lambda)\theta_{t}-\eta\frac{\hat{m_{t}}}{\sqrt{\hat{v}_{t}+\epsilon}}$diamond Update the model parameters with
@@
$\mathbb{E}\left[\left\|\tilde{g_{t}}\right\|^{2}\right]\leq 4(1+q^{2})\left(1+\frac{(1+\omega)^{2}q^{2}}{(1-(1+\omega)q)^{2}}\right)G^{2}.$
@@
$\mathbb{E}\left[\left\|\tilde{g_{t}}\right\|^{2}\right]$ $=\mathbb{E}\left[\left\|\mathcal{C}\left(g_{t}+e_{t}\right)\right\|^{2}\right]$
@@
$\mathbb{E}\left[\left\|\tilde{g_{t}}\right\|^{2}\right]$ $\leq 4(1+q^{2})\left(1+\frac{(1+\omega)^{2}q^{2}}{(1-(1+\omega)q)^{2}}\right)G^{2}.$
@@
$\theta_{t+1}=(1-\eta\lambda)\theta_{t}-\eta_{t}\frac{m^{t}}{\sqrt{v^{t}+\varepsilon}}.$
@@
where $\lambda_{t}=\frac{\lambda}{2}\sum_{i=1}^{t}\left(\frac{1-q}{2}\right)^{i}(t>0),$ $\lambda_{0}=0$ with $10<q<$ 1 and $\|\theta_{t}\|_{\hat{v}_{t}}=\sqrt{\left\langle\theta_{t},\hat{v}_{k}t\otimes\theta_{t}\right\rangle}$. Also let $c_{1}\leq\|\widehat{v}_{t}\|_{\infty}<c_{2}$, then iterates of Algorithm 5 satisfy
@@
Let Assumptions  1 to 5 hold. Let us define $\widetilde{f}\left(\theta_{t}\right)=f(\theta_{t})+\lambda_{t}\|\theta_{t}\|_{\widehat{v}_{t}}^{2}$, where $\lambda_{t}=\frac{\lambda}{2}\sum_{i=1}^{t}\left(\frac{1-q}{2}\right)^{i}(t>0),$ $\lambda_{0}=0$ with $10<q<$ 1 and $\|\theta_{t}\|_{\hat{v}_{t}}=\sqrt{\left\langle\theta_{t},\hat{v}_{k}\otimes\theta_{t}\right\rangle}$. Also define $\Psi_{t}=\widetilde{f}(\theta_{t})+\frac{\eta_{t}}{2c_{1}\beta_{1}}\mathbb{E}\left[\|m_{t}-\nabla f(\theta_{t})\|^{2}\right].$ With $\eta_{t}=\eta\leq\frac{\beta_{1}c_{1}}{2L}\sqrt{\frac{c_{1}}{2c_{2}}}$, Algorithm 5 satisfies
@@ Discussion.
Appendix E Error Feedback applied to GaLore
E.1 Behaviour of the Error Feedback Mechanism

### v2

@@ E.3 Analysis Under PL Condition
$\mathbb{E}[f(\theta_{T+1})]-f^{*}\leq\textstyle\bigl{(}1-\frac{\eta\mu}{C_{0}}\bigr{)}^{T}(f(\theta_{1})-f^{*})+\eta\left(\frac{LC_{0}\sigma^{2}+LC_{0}(C_{1}+C_{2}^{2})G^{2}}{\mu\epsilon}+\frac{(1+C_{1})G^{2}d+C_{1}G^{2}}{\sqrt{\epsilon}}\right)$
$\hfill\textstyle+\,\eta^{2}\left(\frac{3L^{2}C_{0}C^{2}_{1}G^{2}}{2\mu\epsilon^{3/2}}+\frac{(1+2C_{1})C_{1}LG^{2}d}{\epsilon}+\frac{LC_{1}^{2}G^{2}}{2\epsilon}\right).$
@@
To get the rate mentioned in the main part of the paper, we plug in the expression $\eta=\min\{\frac{\epsilon}{4LC_{0}},\frac{2C_{0}\log T}{\mu T}\}$ and collect higher order terms.
$\mathbb{E}[f(\theta_{T+1})]-f^{*}\leq$ $\max\left\{\frac{1}{T^{2}},\left(1-\frac{\epsilon\mu}{4L}\right)^{T}\right\}(f(\theta_{1})-f^{*})$
$+\frac{\log T}{T}\frac{2C_{0}}{\mu}\left(\frac{LC_{0}\sigma^{2}}{\mu\epsilon}+\frac{LC_{0}(C_{1}+C_{2}^{2})G^{2}}{\mu\epsilon}+\frac{(1+C_{1})G^{2}d}{\sqrt{\epsilon}}+\frac{C_{1}G^{2}}{\sqrt{\epsilon}}\right)$
$+\frac{\log^{2}T}{T^{2}}\frac{4C_{0}^{2}}{\mu^{2}}\left(\frac{3L^{2}C_{0}C^{2}_{1}G^{2}}{2\mu\epsilon^{3/2}}+\frac{(1+C_{1})C_{1}LG^{2}d}{\epsilon}+\frac{C_{1}^{2}LG^{2}d}{\epsilon}+\frac{LC_{1}^{2}G^{2}}{2\epsilon}\right)$
$\leq\frac{2\log T}{T}\left(\frac{LC_{0}^{2}}{\mu}\frac{\sigma^{2}+(C_{1}+C_{2}^{2})G^{2}}{\mu\epsilon}+\frac{C_{0}(1+C_{1})(1+d)G^{2}}{\mu\sqrt{\epsilon}}\right)+\widetilde{\mathcal{O}}\left(\frac{G^{4}(G+d)}{T^{2}}\right).$
E.4 Non-convex Analysis with Weight Decay
@@
8:     $\theta_{t+1}=(1-\eta_{t}\lambda)\theta_{t}-\eta\frac{{\color[rgb]{0,0,0}m_{t}}}{\sqrt{{\color[rgb]{0,0,0}v_{t}}+\epsilon}}$diamond Update the model parameters with weight decay
@@
$\mathbb{E}\left[\left\|{\color[rgb]{0,0,0}\tilde{g}_{t}}\right\|^{2}\right]\leq 4(1+q^{2})\left(1+\frac{(1+\omega)^{2}q^{2}}{(1-(1+\omega)q)^{2}}\right)G^{2}.$
@@
$\mathbb{E}\left[\left\|{\color[rgb]{0,0,0}\tilde{g}_{t}}\right\|^{2}\right]$ $=\mathbb{E}\left[\left\|\mathcal{C}\left(g_{t}+e_{t}\right)\right\|^{2}\right]$
@@
$\mathbb{E}\left[\left\|{\color[rgb]{0,0,0}\tilde{g}_{t}}\right\|^{2}\right]$ $\leq 4(1+q^{2})\left(1+\frac{(1+\omega)^{2}q^{2}}{(1-(1+\omega)q)^{2}}\right)G^{2}.$
@@
$\theta_{t+1}=(1-\eta\lambda)\theta_{t}-\eta_{t}\frac{{\color[rgb]{0,0,0}m_{t}}}{\sqrt{{\color[rgb]{0,0,0}v_{t}}+\varepsilon}}.$
@@
where $\lambda_{t}=\frac{\lambda}{2}\sum_{i=1}^{t}\left(\frac{1-q}{2}\right)^{i}$ for $t>0$, $\lambda_{0}=0$ with $10<q<$ 1 and $\|\theta_{t}\|_{\hat{v}_{t}}=\sqrt{\left\langle\theta_{t},{\color[rgb]{0,0,0}\hat{v}_{t}}\otimes\theta_{t}\right\rangle}$. Also let $c_{1}\leq\|\widehat{v}_{t}\|_{\infty}<c_{2}$, then iterates of Algorithm 4 satisfy
@@
Let Assumptions  1 to 5 hold. Define $\Psi_{t}=\widetilde{f}(\theta_{t})+\frac{\eta_{t}}{2c_{1}\beta_{1}}\mathbb{E}\left[\|m_{t}-\nabla f(\theta_{t})\|^{2}\right].$ With $\eta_{t}=\eta\leq\frac{\beta_{1}c_{1}}{2L}\sqrt{\frac{c_{1}}{2c_{2}}}$, Algorithm 4 satisfies
@@ Discussion.
Appendix F Error Feedback applied to GaLore
F.1 Behaviour of the Error Feedback Mechanism